OpenStudy (anonymous):

(This is the law of cosines) How do I solve for angle C, when side a=9 side b=12 side c=15.

OpenStudy (psymon):

c^2 = a^2 + b^2 - 2ab(cosC) would be the formula if I remember correctly.

OpenStudy (psymon):

So we just need to plug in values and then be careful with our algebra from there.

OpenStudy (psymon):

So try plugging in a, b, and c, and see what you're left with. Can check your work after that and help you finish theproblem :3

OpenStudy (anonymous):

15^2 = 9^2 + 12^2 - 2(9)(12) cosC Then 225 = 81 + 144 - 216 cosC then 225 = 9cosC then 25 = cosC Now I'm stuck.

OpenStudy (psymon):

They kind of gave you a trick question, but no, not quite. Notice that in this particular problem, c^2 is the same as a^2 + b^2. If you add 81 and 144 on the right side of your equation, youll get 225, which would cancel out the other side of the equation, leaving you with 0 = -216cosC

OpenStudy (anonymous):

So the -2(a)(b) is attached to the cosC???

OpenStudy (psymon):

Yes, a complete multiplication of.

OpenStudy (anonymous):

Okay. So what do I do about the 0=-216cosC

OpenStudy (psymon):

Sorry, the site went verrrrrrry laggy on me.

OpenStudy (psymon):

Well, essentially it's saying that cosC = 0. So you could do inverse cosine or if you happen to just know where cos is at 0, you'll have your answer for C.

OpenStudy (anonymous):

It's one

OpenStudy (anonymous):

The angle measure for a triangle is 1 degree?

OpenStudy (psymon):

Nope, it'd be pi/2, or 90 degrees.

OpenStudy (anonymous):

Wait, so this is right triangle?????

OpenStudy (psymon):

Its a right triangle :P

OpenStudy (psymon):

Thats why I said its kind of a trick question. But yes, its 90 degrees. So you can now get the other two angles using the law of sines.

OpenStudy (anonymous):

Thanks XDDD